OpenStudy (anonymous):

If M(x, y) is the midpoint of AB with A(4, 6) and B(-8, -12), then which Quadrant is M located in? (Points : 1)

OpenStudy (mrnood):

just find the mid X point (halfway between the2 X ordinates) find the mid Y point(halfway between the2 Y ordinates) this gives the (x,y) coordinates of the midpoint You can plot this, or look a the sign of the 2 values to tell which quadrant it is in

OpenStudy (anonymous):

by mid -pnt. formula \[M(\frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2})\] {4+(-8)}/2=-2=x [6+(-12)]/2=-3=y M(-2,-3) so both x& y coordinates are negative hence 4th quadrant